Grade 12th passAlgebra

If a=7+4root3 & ab=1 then what is the value of 1/a^2+1+b^2

Sonali
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
ab = 1
Hence
b = 1/[7 + 4sqrt(3)] = 7 - 4 sqrt(3)
Now
a² = 97 + 56 sqrt(3)
b² = 97 - 56 sqrt (3)
Now you can solve for 
1/a²    + 1/b²